Syldavia is a country that has 200 workers and neighbouring Borduria has 250 workers. In Syldavia, each worker can make 20 pairs of socks or 30 candles in one year. In Borduria, each worker can make 16 pairs of socks or 20 candles in one year. Who has comparative advantage in producing which good and how will they trade with each other?
1
Expert's answer
2018-12-25T09:54:11-0500

Syldavia is a country that has 200 workers and neighbouring Borduria has 250 workers. If in Syldavia, each worker can make 20 pairs of socks or 30 candles in one year, then the whole country can make 200*20 = 4000 pairs of socks or 200*30 = 6000 candles. Borduria can make 250*16 = 4000 pairs of socks or 250*20 = 5000 candles in one year. So, Syldavia has comparative advantage in producing candles and will sell candles and Borduria has comparative advantage in producing socks and will sell socks